# SpatialVLA: Exploring Spatial Representations for Visual-Language-Action Models

### v2

@@ I Introduction
Generalist robot policies that are capable of interacting with the physical environment, adapting to various embodiments, and performing complex tasks have been a long-standing pursuit in robotics [6, 3, 16, 8, 60]. Recent advances in Vision-Language-Action (VLA) models [7, 29, 5, 32] show a promising paradigm in building such generalist policy by fine-tuning the pre-trained Vision-Language Models (VLMs) [1, 51, 48, 36] on diverse robot data [13, 28, 18]. The key to the success of this paradigm lies in adapting the generalization power of VLMs to numerous robot manipulation tasks, as well as specific architectural designs that synergize the VLM backbone and robot action output head. Nonetheless, existing VLA models are primarily confined to 2D observation inputs and lack precise perception and comprehension of the 3D physical world, where humans may imagine an aligned, canonical, even personally specific 3D workspace for their manipulation. Therefore, an essential question for the field now is how to effectively equip the VLA models with the spatial understanding of the 3D physical world?
@@ III-A The SpatialVLA Model Architecture
As illustrated in Figure. 2, for tokenizing continuous translation and rotation movements, we first normalize each action variable into $[-1,1]$ for each robot environment and statistic the translation and rotation movements $\Delta\mathbf{R}=\{\text{roll, pitch, yaw}\},\;\Delta\mathbf{T}=\{\text{$\phi,\theta,r$}\}$ on the whole dataset mixture (see Section. -F), following with a parameterized Gaussian distribution fitting $\mathcal{N}(\mu^{a},\Sigma^{a})$. Then, the continuous actions are split into M intervals $\mathbf{G}_{i=1,..,\text{M}}=\{[a_{1}=-1,a_{2}),...,[a_{\text{M}},a_{\text{M}}=1]\}$ with equal probability $1/\text{M}$ on each normalized action variable, i.e.,
@@ IV Experiment
Figure 3: Experiment Setup. We evaluate SpatialVLA across 7 robot learning scenarios, 16 real-robot tasks, and 48 simulation setups, focusing on three key aspects: zero-shot control, spatial understanding, and adaptability to new setups. We also conduct a thorough ablation study on a mixed Fractal and Bridge dataset to verify our design decisions.
To answer these questions, as shown in Figure. 3, we evaluate SpatialVLA’s capabilities across a representative spectrum of 7 different robot learning scenarios with 24 real-robot tasks and 3 simulation environments. Firstly, we evaluate SpatialVLA in both SimplerEnv [34] simulation and the real-world WidowX robot platform (BridgeV2 [59] [59] setups), testing its out-of-the-box control capabilities on different robots with setups matching the pre-training dataset. Second, we design 4 special tasks that require precise spatial understanding in 2 different real-world robot environments to test the effectiveness of spatial representations of SpatialVLA. Then, we assess the fine-tuning efficacy of our method in both simulation and real-world settings, including LIBERO [35] and new Franka robot setups, to adapt to new robot environments and tasks. Finally, we conduct comprehensive ablation studies on a mixture of Fractal [6] and BridgeV2 [59] datasets to verify our design decisions in SpatialVLA. For more details on evaluation setups, see Section. -D.
@@ IV-A Performing Zero-shot Robot Control
IV-B Evaluating Spatial Understanding Capability
IV-C Adapting to New Robot Setups

### v4

# Delin Qu∗1, Haoming Song∗1, Qizhi Chen∗1, Yuanqi Yao1, Xinyi Ye1, Yan Ding1, Zhigang Wang1 Jiayuan Gu3, Bin Zhao1, Dong Wang1, Xuelong Li12 1Shanghai AI Laboratory, 2TeleAI, 3ShanghaiTech

@@ I Introduction
Generalist robot policies that are capable of interacting with the physical environment, adapting to various embodiments, and performing complex tasks have been a long-standing pursuit in robotics [6, 3, 16, 8, 60]. Recent advances in Vision-Language-Action (VLA) models [7, 29, 5, 32] show a promising paradigm in building such generalist policy by fine-tuning the pre-trained Vision-Language Models (VLMs) [1, 51, 48, 36] on diverse robot data [13, 28, 18]. The key to the success of this paradigm lies in adapting the generalization power of VLMs to numerous robot manipulation tasks, as well as specific architectural designs that synergize the VLM backbone and robot action output head. Nonetheless, existing VLA models are primarily confined to 2D observation inputs and lack precise perception and comprehension of the 3D physical world — where humans instinctively construct rich, structured mental representations of space, effortlessly aligning objects within a canonical, intuitive, and even personally tailored workspace for manipulation. Therefore, an essential question for the field now is how to effectively equip the VLA models with a profound spatial understanding of the 3D physical world?
@@ III-A The SpatialVLA Model Architecture
As illustrated in Figure. 2, for tokenizing continuous translation and rotation movements, we first normalize each action variable into $[-1,1]$ for each robot environment and statistic the translation and rotation movements $\Delta\mathbf{R}=\{\text{roll, pitch, yaw}\},\;\Delta\mathbf{T}=\{\text{$\phi,\theta,r$}\}$ on the whole dataset mixture (see Section. -F), following with a parameterized Gaussian distribution fitting $\mathcal{N}(\mu^{a},\Sigma^{a})$. Then, the continuous actions are split into M intervals $\mathbf{G}_{i=1,..,\text{M}}=\{[a_{1}=-1,a_{2}),...,[a_{\text{M-1}},a_{\text{M}}=1]\}$ with equal probability $1/\text{M}$ on each normalized action variable, i.e.,
@@ IV Experiment
Figure 3: Experiment Setup. We evaluate SpatialVLA across 7 robot learning scenarios, 16 real-robot tasks, and 48 simulation setups, focusing on three key aspects: zero-shot control, adaptability to new setups, and spatial understanding. We also conduct a thorough ablation study on a mixed Fractal and Bridge dataset to verify our design decisions.
To answer these questions, as shown in Figure. 3, we evaluate SpatialVLA’s capabilities across a representative spectrum of 7 different robot learning scenarios with 24 real-robot tasks and 3 simulation environments. Firstly, we evaluate SpatialVLA in both SimplerEnv [34] simulation and the real-world WidowX robot platform (BridgeV2 [59] [59] setups), testing its out-of-the-box control capabilities on different robots with setups matching the pre-training dataset. Second, we assess the fine-tuning efficacy of our method in both simulation and real-world settings, including LIBERO [35] and new Franka robot setups, to adapt to new robot environments and tasks. Then, we design 4 special tasks that require precise spatial understanding in 2 different real-world robot environments to test the effectiveness of spatial representations of SpatialVLA. Finally, we conduct comprehensive ablation studies on a mixture of Fractal [6] and BridgeV2 [59] datasets to verify our design decisions in SpatialVLA. For more details on evaluation setups, see Section. -D.
@@ IV-A Performing Zero-shot Robot Control
IV-B Adapting to New Robot Setups